# Learning to Generate Object Interactions with Physics-Guided Video Diffusion

### v1

@@ 4 Experiments
We now present our experiments and describe the experimental setup in Section 4.1. Section 4.2 next presents our main results and compares KineMask to the state of the art. Finally Section 4.3 presents a comprehensive analysis with ablations for different components of our method. Further ablations are in Appendix A.3.
@@ 4.1 Setup
We generate two datasets for training and evaluation. Following Section 3.3, we render cubes and cylinders with random colors, moving on textured backgrounds from AmbientCG (AmbientCG, 2018-2025). The first Interactions dataset contains objects moving in random directions and interacting with each other. We also construct a Simple Motion dataset, where isolated objects are moving in random directions without collisions. For both, we generate 10,000 training and 100 test samples. Test videos use disjoint colors and textures compared to training to ensure diversity. We visualize samples of the generated data in Appendix A.2. We also include a set of 50 images, collected from the web or generated with ChatGPT’s image generator (Hurst et al., 2024), used to assess generalization to real scenes with complex objects. Note that unlike Simple Motion and Interactions, does not include ground-truth motion trajectories. We use Tarsier (Wang et al., 2025a) to extract predictive captions $c$.
@@
We adopt CogVideoX-I2V-5B (CogVideoX) as backbone for KineMask. We use ControlNet on the first 8 layers of the model, with weight at 0.5 during training. For $\phi^{\prime\prime}$, we also apply a non-uniform sampling strategy for $f^{*}$, where frame selection is biased toward earlier frames, as rigid-body interactions occur most frequently at the beginning of the simulated sequences. Further details are provided in Appendix A.2. We generate 49 frames at inference.
@@
We consider two baselines using pretrained image-to-video models: CogVideoX (Yang et al., 2025) and Wan2.2-I2V-5B (Wan et al., 2025) (Wan). We prompt both baselines with the same $c_{\text{infer}}$ as used for KineMask. Since we use CogVideoX as a backbone for KineMask, these two methods only differ by our training procedure. Additionally, we evaluate Force Prompting (FP) (Gillman et al., 2025), by mapping our input velocity to input force prompts (Gillman et al., 2025). Force Prompting is also built on CogVideoX. More details on baselines are in Appendix A.2.
@@
For visual quality, we report the Fréchet Video Distance (FVD) (Unterthiner et al., 2019) and the mean squared error (MSE) between generated and ground-truth videos in our synthetic test sets. For motion, we compute the Fréchet Video Motion Distance (FVMD) (Liu et al., 2024a), which isolates motion quality from appearance. Finally, we use SAM2 (Ravi et al., 2025) to extract semantic masks of objects in both generated and ground-truth videos, and compute Intersection over Union (IoU) between them to assess geometric consistency.
Figure 4: User study. We significantly outperform baselines on motion fidelity, interaction quality, and overall physical consistency.
@@ 4.2 Comparison with baselines
We demonstrate the improvements of KineMask against the CogVideoX backbone in Figure 3. We also provide extensive video comparison on all against all baselines in the supplementary material (details are in Appendix A.1). We noticed that CogVideoX suffers mainly from hallucinations, making multiple objects fly or completely disappear. Instead, KineMask generates realistic interactions with other objects if they are present in the path of motion of the initial moving object showing a correct understanding of rigid body dynamics. We also show (Figure 3, bottom left) complex interactions that require implicit 3D understanding, making the glass of juice fall and crash as a result of motion, and multi-object motion and interactions (bottom right). We preserve the input motion direction and object consistency in different types of real-world scenarios, showing a strong generalization of the knowledge acquired from simulated videos.
@@
In Figure 4, we show results of a user study with 30 participants, who perform pairwise comparisons of videos generated by KineMask and baselines for . Participants are presented with initial images and desired object velocities indicated by arrows. We ask participants to evaluate output videos derived from on three different axes: motion fidelity to the control signal (Motion), realism of the object interactions (Interactions), overall physical consistency (Physical). Each answer is collected with a three-answer forced choice format, where we compare our outputs with baselines and we ask for user preferences, allowing also to reply “both have the same quality”. The full details are presented in Appendix A.2. The collected user preferences indicate that our method significantly outperforms all baselines in all questions. This clearly demonstrates the superiority of KineMask in rendering realistic objects interactions.
@@ 4.3.1 Low-level motion control
Direction Speed Object
(a) Degrees of freedom.
Simple Motion
[…]
CogVideoX 158.3 601.1 1504.6 0.051
$\phi^{\prime\prime}$ only 86.3 288.8 201.0 0.237
Ours ($\phi^{\prime}+\phi^{\prime\prime}$) 47.2 160.3 199.8 0.367
$\phi^{\prime}$ w/ masks 24.9 89.7 165.6 0.684
(b) Quantitative results.
Figure 5: Analysis on low-level motion control. In Figure (5(a)), we show different controls of KineMask. We can choose different directions, speed, and objects, opening potential for world modeling. In Table (5(b)), we show that our two-stage training strategy allows to considerably boost results and approach the first-stage training exploiting privileged information ($\phi^{\prime}$ w/ masks).
@@
In Figure 5(a), we present our results on different degrees of freedom of KineMask. Despite being trained on basic synthetic data, KineMask generalizes motion control to complex real-world scenes, coherently with the findings in Gillman et al. (2025). In particular, we show that KineMask achieves disentangled control over different directions, speed, and objects. This allows for a fine-grained evaluation of motion dynamics, precious for world modeling.
@@
We now show the importance of our two-stage training introduced in Section 3.2. We test KineMask on Simple Motion only, to study the effects of motion control in isolation. We first report results of CogVideoX prompted with $c$ as a backbone lower bound. Then, we show the performance of KineMask training directly $\phi^{\prime\prime}$, without first-stage pretraining. As an upper bound, we also evaluate $\phi^{\prime}$ using ground truth object masks provided for every frame at test time. In Table 5(b), we show how our two-stage strategy boosts considerably all metrics compared to training $\phi^{\prime\prime}$ directly, approaching considerably the upper bound $\phi^{\prime}$. We noticed that the IoU is very sensitive to minor displacement errors of the moving objects (0.367 vs 0.684), that do not impact the overall quality of motion as captured by the other metrics.
@@ 4.3.2 Impact of data
We next explore KineMask for generating realistic interactions of rigid bodies. We compare models trained on synthetic samples with and without interactions, and draw conclusions on the capability of VDMs to learn motion. We still use $c_{\emptyset}$ at both training and inference, as in Section 4.3.1.
@@
We evaluate KineMask’s capability to synthesize object interactions depending on the training data. We first test the model trained on Simple Motion and apply it to data. We deliberately set input velocities to directions that should generate collisions among objects in the scene (Figure 6, top). The VDM fails to render realistic collisions. However, KineMask trained on the Interactions set produces accurate interactions for the same input velocities (Figure  6, bottom). Hence, KineMask trained on appropriate data allows to render complex object interactions.
Figure 7: Causal effects. The VDM with KineMask correctly identifies causality, with objects having different effects on the scene depending on their initial velocity.
@@
Moreover, this experiment yields important insights: while VDMs are robust to visual distribution shift, as even synthetic training data do not influence the realism of generated scenes, training on synthetic data with specific motion appears to instead limit the model’s generative capabilities. Indeed, the model trained on Simple Motion is unable to generate interactions. This raises questions on VDMs catastrophic forgetting.
@@
Figure 7 shows results of KineMask trained on Interactions and tested with three different velocities on a scene with interacting objects. As velocity increases, the resulting interactions also change, indicating that the model captures the causal structure of motion. In particular, the final position of the second television varies with the velocity of the first, moving further if the first hits it at higher speeds. This property is valuable for world modeling, as it enables the analysis of different outcomes of object interactions and supports informed planning.
Quantitative evaluation.
We further validate our findings with a quantitative evaluation on Interactions data, following the same protocol as in Section 4.3.1. As shown in Table 1, KineMask achieves the best results when trained on Interactions. Notably, even when trained only on Simple Motion, it still outperforms all baselines, highlighting that KineMask is the most suitable method for rendering motion give an initial velocity.
@@ 4.3.3 High-level text conditioning
Figure 8: Impact of text. Training with rich captions $c$ allows KineMask to generate effects that go beyond the synthetic data used for training, exploiting the prior knowledge of the VDM. Indeed, while we prompt both models with $c_{\text{infer}}$, KineMask trained with $c$ (bottom row) is still able to generate complex effects of the interactions, such as crashing objects, and water effects. See the full prompts in Appendix A.2.
@@
Table 2: Impact of text. Training with $c$, we improve object consistency and general quality on synthetic data.
@@
We also evaluate metrics with different configurations of textual conditioning in Table 2. We consider trainings with $c_{\emptyset}$, using $c_{\emptyset}$ or $c_{\text{infer}}$ at inference. Results demonstrate that using $c$ at training time boosts object consistency (IoU 0.376 vs 0.356 of the second best configuration) and overall realism (FVD 231.3 vs 238.8, MSE on par), while we lose some motion consistency (FMVD 174.4 vs 143.8 of the best model). However, note that Table 2 includes only evaluation on synthetic data, where we have access to a motion ground truth. Considering the improved qualitatives on and the benefit on complex interactions (see Figure 8) we believe the use of captions $c$ during training is still beneficial for the final model to preserve its capabilities.
@@ 5 Discussion
We introduced KineMask, a physics-guided framework that combines low-level motion control with high-level text conditioning to enable video generation of object interactions and effects in complex scenes. Our experiments show that KineMask achieves significant improvements over state-of-the-art models of comparable size, synthesizing realistic multi-object interactions while allowing control over variable object velocities. We believe this methodology can inform future work on world models, with potential implications for robotic manipulation, planning, and other applications of embodied decision making.
While effective, KineMask’s low-level conditioning is limited to velocity, whereas real-world motion also depends on factors such as friction, shape, mass, and air resistance. Incorporating such controls is a promising direction for making VDM-generated motion more physically accurate. In our inference pipeline on real images (Section 3.3), we further use ChatGPT to generate textual descriptions of rigid-body interactions. As shown in Section 4.3.3, combining text-based conditioning with KineMask enhances realism, highlighting the complementary roles of high-level textual guidance and low-level video control. These results support the joint use of both modalities for physically grounded world modeling, and we speculate that advances in multimodal language models (Shukor et al., 2025; Team et al., 2023) may enable text-based physical reasoning to complement video generation.
Acknowledgements
The authors thank Ramazan Fazylov and Viacheslav Iablochnikov for their valuable help and feedback in this work. We also thank to MBZUAI and the Computer Vision Department for their support.
Reproducibility Statement
We currently provide all the details, related to architecture, training stages and hyperparameters in the main paper, Section 4.1, and Appendix A.2. We also include the exact prompts used for the LLMs in Appendix A.2. To further improve reproducibility, will release data, the model, and the source code, along with usage instructions.
@@ A.1 Video qualitative results
For a comprehensive qualitative evaluation on videos, we refer the reader to the attached supplementary material, in which we provide qualitative examples for all images animated with KineMask and baselines. We organize the results by using a website that can be easily opened in any browser. To visualize the videos, simply click on the index.html file after unzipping the material. Please note that for transparency we include randomly selected outputs from all images in . Hence, quality across samples may differ.
@@ A.2 Additional details
We report here the prompt used for $c$ generation at training time and $c_{\text{infer}}$ used for inference time. We prompt Tarsier (Wang et al., 2025a) to extract $c$ with: “Describe the video in a concise way, covering all motions and collisions between objects.”. This simple prompt is already resulting in sufficiently good descriptions. For inference, instead, we rely on GPT prompting with in-context learning examples to extract suitable descriptions. We report the prompt below, along the associated in-context examples in Figure 9.
@@
Inference-time $c_{\text{infer}}$ generation prompt Reference: The first five attached examples are the first frames of simulated videos that contain multiple objects on a surface, in each example one or two objects can start moving, and these objects can or cannot collide with other objects depending on their initial velocity and direction. The range of velocities for the initial motion of these objects is from 0.5m/s to 1.5m/s. - In the first initial frame the green cube moves at 1.23m/s and the red cylinder at 1.07m/s. The entire video description for this video is: ”On a wooden floor, there are four objects: a red cylinder, a green cube, a yellow cube, and a white cylinder. The red cylinder starts in the center and moves to the right, while the green cube moves to the left. The red cylinder continues to move to the right and eventually collides slightly with the white cylinder. The green cube continues to move left and collides slightly with the yellow cube. The red cylinder and the white cylinder remain stationary after the collision.”. - In the second initial frame the white cube moves at 0.67m/s and the purple cube at 1.35m/s, the entire description for this video is: ”On a grassy background, there are a yellow cylinder, a purple cube, a pink cube, and a white cube. The white cube moves towards the yellow cylinder and collides slightly with it. The white cube then stops next to the yellow cylinder. The purple cube moves towards the pink cube and collides with it, moving it towards the top right, after the collision all objects remain stationary.”. - In the third initial frame the white cube moves at 1.5m/s and the light blue cube at 1.43m/s, the entire description for this video is: ”Four cubes of different colors (yellow, green, blue, and purple) are on a textured, brown surface. The blue cube moves towards the green cube, colliding with it and pushing it downwards. The yellow cube moves towards the pink cube, they collide and the purple cube moves out of the frame. After the collision the blue and green cylinder are static next to each other, while the yellow cube is positioned in the top right part of the frame.”. - In the fourth initial frame, the gray cube moves at 0.64m/s, the entire description for this video is: ”The scene is set on a brick ground with patches of green moss. A white cylinder is stationary in the background. A gray cube, initially positioned on the left side of the frame, moves horizontally to the right and eventually stops in the center of the frame. Two pink cubes, one on the right side and another slightly behind it, remain stationary throughout the sequence.”. - In the fifth initial frame the white cylinder moves at 0.71m/s and the pink cylinder moves at 0.55m/s. The entire description for this video is: ”The scene consists of a grassy background with four 3D shapes: a purple cube, a white cylinder, a pink cylinder, and another purple cube. The white cylinder and the pink cylinder move slightly to the left. The purple cubes in the foreground remain stationary throughout the sequence.”. Objective: -The last attached initial frame consists of a real picture, in this case the xxx on the xxx moves to the xxx at xxxm/s !!. According to this information and taking into account the provided examples from simulated videos, please provide an entire description, predicting what is going to happen in this scene. The provided video description should follow the same style and words used in the given simulated examples !! , reason about the real scene and describe if collisions happen or not, if they do, describe them, consider the distance and velocity between objects in the real scene to determine how the objects move, and how strong the collisions are. Make sure the description would include realistic effects if they happen produced by the movement or collisions that may occur between objects, for example movement of liquids, steam effects, falling objects, etc. If these effects are not present in the scene please do not include anything related to them in the prompt !!. Do not include velocity values in it, friction effects, information from the surface, inclinations, rotations, or produced sounds.
@@
We train KineMask using bf16 mixed precision for 1000 steps with a total batch size of 40 and saved checkpoints every 250 steps. Our backbone is initialized from the pretrained THUDM/CogVideoX-5b-I2V model. For the Controlnet we used the first 8 transformer layers, a downscaling factor of 8 and a control-weight of 0.5. We use AdamW as the optimizer with a learning rate of $1\times 10^{-4}$, $\beta_{1}$=0.9, $\beta_{2}$=0.95, a cosine-with-restart learning-rate schedule and a max gradient norm of 1.0.
@@
We compare KineMask with Force Prompting, CogvideoX, and Wan. All methods use an input prompt and an initial image. Additionally, Force Prompting takes the point coordinates of where to apply the force, the force prompt, and the angle of movement. In all cases, we apply the same prompt $c_{\text{infer}}$ that we extract by querying GPT. In the extracted prompts, the moving object and its interactions are always described. Our results demonstrate that for low-level control (CogvideoX and Wan), the simple application of textual conditioning results in ambiguous information, ultimately yielding generation artifacts.
@@
We report here the prompts used for rendering Figure 8, that we omitted due to the lack of space.

### v2

@@ 4 Experiments
We describe the experimental setup in Section 4.1. Section 4.2 presents our main results and compares KineMask to the state of the art. Finally, Section 4.3 shows a comprehensive analysis with ablations for different components of our method. Further ablations are in Appendix A.3.
@@ 4.1 Setup
We generate two datasets for training and evaluation. Following Section 3.3, we render cubes and cylinders with random colors, moving on textured backgrounds from AmbientCG [6]. The first Interactions dataset contains objects moving in random directions and interacting with each other. We also construct a Simple Motion dataset, where isolated objects are moving in random directions without collisions. For both, we generate 10,000 training and 100 test samples. Test videos use disjoint colors and textures compared to training to ensure diversity. We visualize samples of the generated data in Appendix A.2. We also include a set of 50 images, collected from the web or generated with ChatGPT’s image generator [29], used to assess generalization to real scenes with complex objects. Note that unlike Simple Motion and Interactions, does not include ground-truth motion trajectories. We use Tarsier [58] to extract predictive captions $c$.
Direction Speed Object
Figure 5: Degrees of freedom. We show control of different aspects of KineMask outputs. We can choose different directions (left), speed (middle), and objects to move (right), opening potential for world modeling.
@@
We adopt CogVideoX-I2V-5B (CogVideoX) as backbone for KineMask. We use ControlNet on the first 8 layers of the model, with weight at 0.5. For $\phi^{\prime\prime}$, we also apply a non-uniform sampling strategy for $f^{*}$, where frame selection is biased toward earlier frames, as rigid body interactions occur most frequently at the beginning of the simulated sequences. Further details are provided in Appendix A.2. We generate 49 frames at inference.
@@
We consider two baselines using pretrained image-to-video models: CogVideoX [61] and Wan2.2-I2V-5B [55] (Wan). We prompt both baselines with the same $c_{\text{infer}}$ as used for KineMask. Since we use CogVideoX as a backbone for KineMask, these two methods only differ by our training procedure. We choose TORA [66], also based on CogVideoX, as a drag-based baseline. Since TORA requires the full drag trajectory, we draw it starting from the target object depending on the input velocity (see Appendix A.2). We also include MotionI2V [49] as mask-constrained drag method, that we prompt using SAMv2 masks and a drag trajectory similar to TORA. Finally, we evaluate Force Prompting (FP) [20], by mapping our input velocity to input force prompts [20]. Force Prompting is also built on CogVideoX. More details are in Appendix A.2.
@@
For visual quality, we report the Fréchet Video Distance (FVD) [53] and the mean squared error (MSE) between generated and ground-truth videos in our synthetic test sets. For motion, we compute the Fréchet Video Motion Distance (FVMD) [38], which isolates motion quality from appearance. Finally, we use SAM2 [46] to extract semantic masks of objects in generated videos, and compute Intersection over Union (IoU) with ground-truth masks and assess geometric consistency.
Simple Motion
[…]
CogVideoX 158.3 601.1 1504.6 0.051
2nd stage only 86.3 288.8 201.0 0.237
Ours (1st + 2nd stage) 47.2 160.3 199.8 0.367
1st stage w/ masks 24.9 89.7 165.6 0.684
Table 1: Ablation on training. We show that our two-stage training strategy considerably boosts results and approaches an upper bound exploiting privileged information (1st stage w/ masks).
@@ 4.2 Comparison with baselines
We demonstrate qualitative improvements on scenes. While for space reasons we compare only with the backbone CogVideoX in Figure 3, we include detailed comparisons against all baselines in the Appendix. We noticed that CogVideoX suffers from unrealistic interactions, making objects fly or disappear. Instead, KineMask generates realistic interactions with other objects if they are present in the path of motion of the initial moving object, showing a correct understanding of rigid body dynamics. We also show (Figure 3, bottom left) complex interactions that require implicit 3D understanding, making the glass of juice fall and crash as a result of motion, and multi-object motion and interactions (bottom right). Our realistic real world outputs show a strong generalization of the knowledge acquired from simulated videos.
@@
In Figure 4, we show results of a user study with 30 participants, who perform pairwise comparisons of videos generated by KineMask and baselines on . Participants are presented with initial images and input object velocities indicated by arrows. We ask them to evaluate outputs on three different axes: motion fidelity to the control signal (Motion), realism of the object interactions (Interactions), overall physical consistency (Physical). Each answer is collected with a three-answer forced choice format, where we compare our outputs with baselines and we ask for user preferences, allowing also to reply “both have the same quality”. The full details are presented in Appendix A.2. The collected user preferences indicate that our method significantly outperforms all baselines in all questions. This demonstrates that existing methods are unsuitable for rendering interactions with starting conditions only. Kinemask, instead, renders realistic interactions.
@@ 4.3.1 Low-level motion control
In Figure 5, we present results on different degrees of freedom of KineMask. Despite being trained on basic synthetic data, KineMask generalizes motion control to complex real-world scenes, coherently with the findings in [20]. In particular, we show that KineMask achieves disentangled control over different directions, speed, and objects. This allows for a fine-grained evaluation of motion dynamics, precious for world modeling.
@@
In Table 1 we show the importance of our two-stage training introduced in Section 3.2. We test KineMask on Simple Motion only, to study the effects of motion control in isolation. We first report results of CogVideoX prompted with $c$ as a lower bound. As additional baseline, we train directly $\phi^{\prime\prime}$ (2nd stage only), without first-stage pretraining. As an upper bound, we also evaluate $\phi^{\prime}$ using ground truth object masks provided for every frame at test time (1st stage w/ masks). Our two-stage strategy boosts considerably all metrics compared to training $\phi^{\prime\prime}$ directly, approaching the upper bound. We noticed that the IoU is very sensitive to minor displacement errors of the moving objects (0.367 vs 0.684), that do not impact the overall quality of motion as captured by the other metrics.
@@ 4.3.2 Impact of data
Next, we explore the effects of training data on KineMask for generating realistic rigid body interactions. We compare models trained on synthetic samples with and without interactions, and draw conclusions on the capability of VDMs to learn motion. We still use $c_{\emptyset}$ at both training and inference, as in Section 4.3.1.
@@
We evaluate KineMask’s capability to synthesize object interactions depending on the training data. We first test the model trained on Simple Motion and apply it to data. We deliberately set input velocities to directions that should generate collisions among objects in the scene. The VDM fails to render realistic collisions (Figure 6, top). However, KineMask trained on the Interactions set produces accurate interactions for the same input velocities (Figure  6, bottom). Hence, KineMask trained on appropriate data allows to render complex object interactions. Moreover, this experiment yields important insights: while VDMs are robust to visual distribution shift, as even synthetic training data do not influence the realism of generated scenes, their preservation of motion understanding seems to be more impacted, since training on Simple Motion prevents to generate interactions. This raises questions on VDMs catastrophic forgetting.
Figure 7: Causality effects. On top, we show how different velocity impacts motion effects, suggesting causal understanding. At the bottom, we show that the environment impacts rendering results for the same input velocity but different objects, suggesting mass awareness.
Quantitative evaluation.
We validate the importance of specific data with a quantitative evaluation on Interactions, following the same protocol as in Section 4.3.1. As shown in Table 2, KineMask achieves the best results when trained on Interactions. Notably, even when trained only on Simple Motion, it greatly outperforms CogVideoX, highlighting the need of our research.
@@
Figure 7 shows results of KineMask trained on Interactions and tested with different inputs on a scene with interacting objects. On top, as velocity increases, the interactions also change, indicating that the model captures causality of collisions. At the bottom, with the same velocity, we inpaint the input image, removing the interacting object, or replacing it with objects of different mass. The different final position of the moving object demonstrates understanding of mass in collisions. Note that this result is challenging for drag-based methods since they imply prior knowledge of the object final position. KineMask instead maps different target positions to the same input, depending on the environment.
@@ 4.3.3 High-level text conditioning
Figure 8: Impact of text. Training with rich captions $c$ makes KineMask generate effects that go beyond the synthetic data used for training, exploiting the prior knowledge of the VDM. Indeed, while we prompt both models with $c_{\text{infer}}$, KineMask trained with $c$ (bottom row) is able to generate complex effects of the interactions, such as crashing objects, and water effects. See the full prompts in Appendix A.2.
@@
Table 3: Ablation on captions. Training with $c$, we improve object consistency and general quality on synthetic data.
@@
We also evaluate metrics with different textual conditioning in Table 3. We consider trainings with $c_{\emptyset}$, using $c_{\emptyset}$ or $c_{\text{infer}}$ at inference. Results demonstrate that using $c$ at training time boosts object consistency (IoU 0.376 vs 0.356 of the second best configuration) and overall realism (FVD 231.3 vs 238.8, MSE on par), while we lose some motion consistency (FMVD 174.4 vs 143.8 of the best model). However, note that Table 3 includes only evaluation on synthetic data, where we have access to a motion ground truth. Considering the improved qualitatives on and the benefit on complex interactions (see Figure 8) we believe the use of captions $c$ during training is still beneficial for the final model to improve robustness and preserve its capabilities.
@@ 5 Discussion
We introduced KineMask, a framework to enable video generation of object interactions and effects in complex scenes. Our experiments show that KineMask achieves significant improvements over state-of-the-art models of comparable size, synthesizing realistic multi-object interactions while allowing control over variable object velocities. We believe this methodology can inform future work on world models, with potential implications for robotic manipulation, planning, and embodied decision making.
Limitations and future work.
While effective, KineMask’s low-level conditioning is limited to velocity, whereas real-world motion also depends on factors such as friction, shape, mass, and air resistance. Incorporating such controls is a promising direction for making VDM-generated motion more physically accurate. Moreover, we believe that extending to soft-body interactions will further improve world modeling capabilities. In our inference pipeline on real images (Section 3.3), we further use ChatGPT to generate textual descriptions of rigid-body interactions. As shown in Section 4.3.3, combining text-based conditioning with KineMask enhances realism, highlighting the complementary roles of high-level textual guidance and low-level video control. These results support the joint use of both modalities for physically grounded world modeling, and we speculate that advances in multimodal language models [51, 52] may enable text-based physical reasoning to complement video generation.
@@ A.1 Video qualitative results
For a comprehensive qualitative evaluation on videos, we refer the reader to the attached supplementary material, in which we provide qualitative examples for all images animated with KineMask and baselines. We organize the results by using a website that can be easily opened in any browser. To visualize the videos, simply click on the KineMask-main/index.html file after unzipping the material. Please note that for transparency we include randomly selected outputs from all images in and common failure cases.
@@ A.2 Additional details
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 10: Qualitative comparison with all baselines.
@@
We report here the prompt used for $c$ generation at training time and $c_{\text{infer}}$ used for inference time. We prompt Tarsier [58] to extract $c$ with: “Describe the video in a concise way, covering all motions and collisions between objects.”. This simple prompt is already resulting in sufficiently good descriptions. For inference, instead, we rely on GPT prompting with in-context learning examples to extract suitable descriptions. We report the prompt below, along the associated in-context examples in Figure 11.
Inference-time $c_{\text{infer}}$ generation prompt Reference: The first five attached examples are the first frames of simulated videos that contain multiple objects on a surface, in each example one or two objects can start moving, and these objects can or cannot collide with other objects depending on their initial velocity and direction. The range of velocities for the initial motion of these objects is from 0.5m/s to 1.5m/s.
- In the first initial frame the green cube moves at 1.23m/s and the red cylinder at 1.07m/s. The entire video description for this video is: ”On a wooden floor, there are four objects: a red cylinder, a green cube, a yellow cube, and a white cylinder. The red cylinder starts in the center and moves to the right, while the green cube moves to the left. The red cylinder continues to move to the right and eventually collides slightly with the white cylinder. The green cube continues to move left and collides slightly with the yellow cube. The red cylinder and the white cylinder remain stationary after the collision.”.
- In the second initial frame the white cube moves at 0.67m/s and the purple cube at 1.35m/s, the entire description for this video is: ”On a grassy background, there are a yellow cylinder, a purple cube, a pink cube, and a white cube. The white cube moves towards the yellow cylinder and collides slightly with it. The white cube then stops next to the yellow cylinder. The purple cube moves towards the pink cube and collides with it, moving it towards the top right, after the collision all objects remain stationary.”.
- In the third initial frame the white cube moves at 1.5m/s and the light blue cube at 1.43m/s, the entire description for this video is: ”Four cubes of different colors (yellow, green, blue, and purple) are on a textured, brown surface. The blue cube moves towards the green cube, colliding with it and pushing it downwards. The yellow cube moves towards the pink cube, they collide and the purple cube moves out of the frame. After the collision the blue and green cylinder are static next to each other, while the yellow cube is positioned in the top right part of the frame.”.
- In the fourth initial frame, the gray cube moves at 0.64m/s, the entire description for this video is: ”The scene is set on a brick ground with patches of green moss. A white cylinder is stationary in the background. A gray cube, initially positioned on the left side of the frame, moves horizontally to the right and eventually stops in the center of the frame. Two pink cubes, one on the right side and another slightly behind it, remain stationary throughout the sequence.”.
- In the fifth initial frame the white cylinder moves at 0.71m/s and the pink cylinder moves at 0.55m/s. The entire description for this video is: ”The scene consists of a grassy background with four 3D shapes: a purple cube, a white cylinder, a pink cylinder, and another purple cube. The white cylinder and the pink cylinder move slightly to the left. The purple cubes in the foreground remain stationary throughout the sequence.”.
@@
Objective: -The last attached initial frame consists of a real picture, in this case the xxx on the xxx moves to the xxx at xxxm/s !!. According to this information and taking into account the provided examples from simulated videos, please provide an entire description, predicting what is going to happen in this scene. The provided video description should follow the same style and words used in the given simulated examples !! , reason about the real scene and describe if collisions happen or not, if they do, describe them, consider the distance and velocity between objects in the real scene to determine how the objects move, and how strong the collisions are. Make sure the description would include realistic effects if they happen produced by the movement or collisions that may occur between objects, for example movement of liquids, steam effects, falling objects, etc. If these effects are not present in the scene please do not include anything related to them in the prompt !!. Do not include velocity values in it, friction effects, information from the surface, inclinations, rotations, or produced sounds.
@@
We train KineMask using bf16 mixed precision for 1000 steps with a total batch size of 40 and saved checkpoints every 250 steps. Our backbone is initialized from the pretrained THUDM/CogVideoX-5b-I2V model. For the Controlnet we used the first 8 transformer layers, a downscaling factor of 8 and a control-weight of 0.5. We use AdamW optimizer with a learning rate of $1\times 10^{-4}$, $\beta_{1}$=0.9, $\beta_{2}$=0.95, a cosine-with-restart learning-rate schedule and a max gradient norm of 1.0.
@@
We compare KineMask with Force Prompting, TORA, Motion-I2V, CogvideoX and Wan. All methods use an input prompt and an initial image. Additionally, Force Prompting takes the point coordinates of where to apply the force, the force prompt, and the angle of movement. Tora and Motion-I2V need the entire motion trajectory, we give the initial and final points. In all cases, we apply the same prompt $c_{\text{infer}}$ that we extract by querying GPT. In the extracted prompts, the moving object and its interactions are always described.
@@
We report here the prompts for rendering Figure 8, that we omitted due to the lack of space.
@@
A.3 Additional results
Test on different model backbones
We tested our proposed approach on two model backbones CogVideoX-5B-I2V and Wan2.2-I2V-5B, and performed a quantitative comparison on the evaluation set of our generated synthetic data, by training and testing on Interactions. Table 4 shows that both KineMasks versions outperform their original zero-shot backbones by large margins respectively. On the other hand, we noticed that KineMask-CogVideoX achieves better performances than its Wan counterpart, which also relates when comparing the performance of the original backbones when these are tested in a zero-shot manner, where CogVideoX achieves better performances than Wan. Figure 12 shows a comparison of real-world generations using KineMasks with both backbones. We noticed that while both backbones generate plausible object interactions, the model that used CogVideoX as backbone showed better object consistency. For example, using KineMask-Wan in Figure 12 top, the candles show some overlapping (Red Square) when the collision is generated, similarly in Figure 12 bottom, the fruit increased its size after the collision. On the other hand, we noticed that using KineMask-CogVideoX we obtained better rigid-body collisions. Due to this, we chose KineMask-CogvideoX (KineMask) as our main model.
Interactions
[…]
Wan Baseline 431.4 977.7 3649.8 0.192
KineMask 211.0 475.6 968.4 0.257
CogVideoX Baseline 344.6 807.3 3514.9 0.192
KineMask 158.7 250.7 143.8 0.355
Table 4: Comparison of different model backbones. We compare two versions of KineMask using different base models.
Figure 12: Comparison of model backbones.
Figure 13: Qualitative comparison of KineMask vs PhysGen3D.
Qualitative comparison against all baselines
Figure 10 presents a complete qualitative comparison between KineMask and all baselines. Beyond the issues observed with CogVideoX (discussed in Section 4.2), Wan exhibits strong hallucinations and unrealistic object interactions. On the other hand, Drag-based methods such as Force-Prompting and Motion-I2V commonly suffer from slow object motion when objects are moved toward others, failing to synthesize proper interactions. Finally, TORA often produces hallucinations and ignores collisions, resulting in objects being placed incorrectly or stacked on top of one another. KineMask successfully generate plausible motion and object collisions according to the user input, showing its capability to correctly understand causality, and dynamics created due to collisions, and showing the generalization of the knowledge obtained from synthetic data in complex real-world scenes, even when some 3D understanding is necessary in order to render collisions with the floor.